OpenStudy (anonymous):

Solve: 5(x + 6) 3x - 4 x > -17 x >17 x <-17 x < -5

OpenStudy (anonymous):

you miss equality sign?

OpenStudy (anonymous):

5(x + 6) > 3x - 4 or 5(x + 6) < 3x - 4

OpenStudy (anonymous):

\(\leq\) or \(\geq\) probably because she copied and those didn't copy usually :d

OpenStudy (anonymous):

OOPS THE SIGN GOES NEXT TO THE 6 \[\ge\]

OpenStudy (anonymous):

Solve: 5(x + 6) \[\le\] 3x - 4